# Quantum Abduction: A New Paradigm for Reasoning under Uncertainty

### v1

@@ 4.2 Literary Demonstration: Murder on the Orient Express
Figure 5: Quantum abductive diagnosis sustains co-activation and permits parallel, low-regret therapy before evidential collapse.
@@ 4.4 Scientific Theory Change
Astrophysics: Dark Matter vs. MOND
@@
For decades, neither model has been definitively falsified. Dark matter has accumulated supporting evidence from cosmology (CMB anisotropies, lensing), but direct detection experiments remain null [22, 11]. MOND captures empirical regularities in galactic dynamics but resists integration into relativistic cosmology. Both thus persist in a suspended explanatory state, as represented in Figure 6, their amplitudes oscillating with new experimental results.
@@
Figure 6: Scientific superposition: competing models co-exist until decisive evidence induces collapse.
Geology: Continental Drift $\rightarrow$ Plate Tectonics
Another striking example of quantum abductive dynamics is the history of continental drift. In 1912, meteorologist Alfred Wegener proposed that continents once formed a single landmass (Pangaea) and had since drifted apart. His evidence included fossil correlations across oceans, paleoclimatic traces (e.g. glacial striations in now-tropical zones), and geological continuities between Africa and South America.
Yet geology at the time was dominated by fixist paradigms: mountains were thought to arise from the Earth’s contraction as it cooled, and continents were assumed immobile. As a result, Wegener’s hypothesis was widely ridiculed, especially because he lacked a plausible mechanism to explain how continents could move.
@@
The wave function did not collapse until the 1960s, when new instruments revealed seafloor spreading, subduction zones, and the dynamics of the lithosphere. With plate tectonics [8], the two alternatives merged into a richer synthesis, represented in Figure 7: Wegener was right about drift, but wrong about the mechanism. His bold intuition thus collapsed into modern geology’s most unifying paradigm.
@@
[…]
Figure 7: Geological superposition: long-duration coexistence of competing models resolved by new mechanisms and measurements.
@@
Case Study Classical Abduction Quantum Abduction Medical Diagnosis (Botulism vs. GBS) Seek single best diagnosis; early resolution; risk of excluding valid alternatives. Maintain co-activation; parallel low-regret therapy; collapse on decisive evidence. Astrophysics (Dark Matter vs. MOND) Commit to dominant paradigm; treat the other as fringe. Keep both entangled; design tests targeting interference zones; defer collapse. Geology (Drift vs. Fixism) Reject Drift for lack of mechanism; persist with Fixism. Sustain superposition until mechanism appears (tectonics); then collapse.
@@ 5.3 Set-covering Approaches
Set-covering abduction, pioneered by Reggia and colleagues [18, 19], frames explanation as a combinatorial optimization task: find minimal sets of causes that “cover” observed effects. Widely applied in medical diagnosis and fault detection, this approach is computationally tractable and intuitively appealing in domains with clear causal mappings. Variants such as Parsimonious Covering Theory further emphasized minimality as a criterion of explanatory adequacy.
@@ 5.4 Integrative Assessment
Taken together, these traditions demonstrate the richness of abductive research in AI, from logical proof systems to probabilistic and diagnostic optimization frameworks. Yet they converge on a common stance: explanations are rivals, and reasoning ends when one prevails. Quantum abduction reconfigures this picture. By embedding hypotheses in semantic Hilbert spaces, allowing them to interfere constructively or destructively, and enabling collapse into composite syntheses, it offers an alternative paradigm—one aligned with how human reasoning sustains ambiguity, navigates contradictions, and generates emergent insights.
We compare these four directions in the formalization and implementation of abduction in Figure 8, emphasizing the pluses and minuses in their ability to unlock the potential of investigative flexibility and explorative creativity intrinsic to this mode of reasoning.
Logic-based Abduction Bayesian Abduction Set-covering Abduction ✓ Rigorous logical scaffolding ✗ Eliminative pruning ✗ Fixed hypothesis space ✗ Contradictions excluded ✓ Handles uncertainty via priors ✗ Fixed hypothesis space ✗ Hypotheses as weighted alternatives ✗ Ultimately eliminative (MPE) ✓ Efficient in diagnosis ✗ Minimality via elimination ✗ Predefined cause–effect mapping ✗ Cannot sustain ambiguity Quantum Abduction ✓ Superposition of hypotheses ✓ Interference (constructive/destructive) ✓ Dynamic synthesis of hybrids ✓ Collapse aligned with evidence
Figure 8: Contrasting abductive frameworks. Classical logics and set-covering are eliminative within fixed spaces; Bayesian abduction handles uncertainty but treats hypotheses as weighted alternatives and still collapses to a single MPE. Quantum abduction sustains superposition, models interference, and collapses to dominant or hybrid explanations.
6 Conclusion and Outlook
This paper has introduced quantum abduction as a novel framework for explanatory reasoning. Unlike classical approaches—logic-based, Bayesian, or set-covering—that treat hypotheses as rivals and eliminate them until one survives, quantum abduction maintains explanatory candidates in superposition, allows them to interfere constructively or destructively, and collapses only when evidence supports a coherent (possibly hybrid) synthesis. Through case studies drawn from forensic investigation, literature, medical diagnostics, and scientific theory change, we illustrated how this perspective provides a richer account of how human reasoning sustains ambiguity, navigates contradictions, and generates emergent explanations. Implementation guidelines sketched how semantic embeddings, projection operators, interference matrices, and amplitude dynamics can be combined into a computational pipeline aligned with this model.
Looking forward, two main directions stand out:
  • Formalization. A complete logical framework for quantum abduction remains to be developed. This may borrow from existing proof-theoretic traditions (e.g. sequent calculi), but will need to extend them with quantum-inspired machinery such as Hilbert spaces, projection operators, and interference dynamics. At the same time, subsymbolic and statistical methodologies—NLP embeddings, semantic similarity models, and generative LLMs—can be integrated to create a hybrid paradigm that bridges symbolic clarity with statistical power.
  • Implementation. Beyond theoretical development, an applied toolkit is needed. Such a system would allow experts to encode hypotheses and observations, visualize interference maps, and receive natural language synthesis of composite explanations. We envision applications in forensic reasoning, clinical diagnostics, and scientific discovery, where complex, contradictory evidence must be integrated into coherent and interpretable accounts.
In sum, quantum abduction reframes explanation as synthesis rather than elimination. By combining the structural elegance of symbolic reasoning with the adaptive strength of subsymbolic methods, it opens the path toward reasoning systems capable of matching the flexibility and creativity of human inquiry.
Funding

### v2

@@ 4.3.3 Clinical Alignment
Figure 3: Quantum abductive diagnosis sustains co-activation and permits parallel, low-regret therapy before evidential collapse. (Solid arrows: evidence projection; dashed arrows: contribution to the entangled state.)
@@ 4.4 Scientific Theory Change
4.4.1 Astrophysics: Dark Matter vs. MOND
@@
For decades, neither model has been definitively falsified. Dark matter has accumulated supporting evidence from cosmology (CMB anisotropies, lensing), but direct detection experiments remain null [12, 13]. MOND captures empirical regularities in galactic dynamics but resists integration into relativistic cosmology. Both thus persist in a suspended explanatory state, as represented in Figure 4a, their amplitudes oscillating with new experimental results.
@@
(a) Astrophysics: Dark Matter vs. MOND
[…]
(b) Geology: Drift vs. Fixism
Figure 4: Scientific superposition in two domains. (a) Astrophysics: dark matter and MOND coexist in an entangled explanatory state until decisive evidence induces collapse. (b) Geology: Wegener’s continental drift and fixist models remain in superposition until the mechanism of plate tectonics emerges, yielding a richer synthesis. (Solid arrows: evidence projection; dashed arrows: contribution to the entangled state.)
4.4.2 Geology: Continental Drift $\rightarrow$ Plate Tectonics
Another striking example of quantum abductive dynamics is the history of continental drift. In 1912, meteorologist Alfred Wegener proposed that continents once formed a single landmass (Pangaea) and had since drifted apart. His evidence included fossil correlations across oceans, paleoclimatic traces (e.g., glacial striations in now-tropical zones), and geological continuities between Africa and South America.
Yet geology at the time was dominated by fixist paradigms: mountains were thought to arise from the Earth’s contraction as it cooled, and continents were assumed immobile. As a result, Wegener’s hypothesis was widely ridiculed, mainly because he lacked a plausible mechanism to explain how continents could move.
@@ 4.4.2 Geology: Continental Drift $\rightarrow$ Plate Tectonics
The wave function did not collapse until the 1960s, when new instruments revealed seafloor spreading, subduction zones, and the dynamics of the lithosphere. With plate tectonics [14], the two alternatives merged into a richer synthesis, represented in Figure 4b: Wegener was right about drift, but wrong about the mechanism. His bold intuition thus collapsed into modern geology’s most unifying paradigm (this example illustrates a central feature of quantum abduction: instead of enforcing early elimination of hypotheses—as classical abductive frameworks often do—the model allows suspended coexistence until decisive evidence becomes available. Historically, continental drift was rejected not because it failed to explain observations, but because it lacked a mechanism; quantum abduction formalizes precisely this long-lived state of justified non-elimination.).
@@
Case Study Classical Abduction Quantum Abduction Medical Diagnosis (Botulism vs. GBS) Seek the single best diagnosis; early resolution; risk of excluding valid alternatives. Maintain co-activation; parallel low-regret therapy; collapse on decisive evidence. Astrophysics (Dark Matter vs. MOND) Commit to the dominant paradigm; treat the other as fringe. Keep both entangled; design tests targeting interference zones; defer collapse. Geology (Drift vs. Fixism) Reject Drift for lack of mechanism; persist with Fixism. Sustain superposition until the mechanism appears (tectonics); then collapse.
4.6 Key Takeaway: The Usefulness of Pre-Collapse
In high-stakes settings, the pre-collapse state is not a temporary inconvenience but a valuable analytical product. While the case studies in the previous section focus on law, medicine, and scientific theory change, the same dynamics extend naturally to domains such as intelligence analysis and emergency management. In every setting, interference patterns reveal where evidence fails to discriminate, and amplitude trends indicate when to act, what to collect next, and how to document uncertainty with transparency.
4.6.1 Law (Reasonable Doubt)
A distribution such as $|\alpha_{H_{1}}|^{2}=0.55$ (guilt), $|\alpha_{H_{2}}|^{2}=0.35$ (alternative), and an emergent $H_{3}=0.10$ identifies evidential gaps, clarifies burdens, supports proportional or qualified verdicts, and provides a structured trajectory for appellate review (cf. Bossetti, Section 4.1).
4.6.2 Intelligence (Collection Planning)
Amplitudes over $H_{1}$ (centralized), $H_{2}$ (autonomous), and $H_{3}$ (hybrid) guide resource allocation proportionally to $|\alpha_{i}|^{2}$. Positive $I_{12}$ highlights the need for discriminating evidence (e.g., communication topology), while growth in $\alpha_{H_{3}}$ suggests preparing hybrid-network models.
4.6.3 Medicine (Parallel Treatment)
When botulism and GBS/MFS maintain high but mutually destructive amplitudes, pre-collapse supports low-regret dual therapy, provides an ethical and administrative rationale for resource use, and indicates when commitment becomes safe as amplitudes diverge (see Figure 3).
4.6.4 Science (Avoid Premature Closure)
Amplitude tracking for drift vs. fixism (historically) or dark matter vs. MOND (today) highlights which evidence classes are decisive and which are merely reinforcing, converting rivalry into targeted exploration rather than premature paradigm commitment.
4.6.5 Crisis Response (Public Safety)
Under ambiguous causes, amplitudes justify graduated protocols—e.g., simultaneous HAZMAT deployment, security lockdown, and meteorological monitoring—and provide a clear after-action trail explaining branch prioritization.
4.6.6 Institutional Value
Pre-collapse visibility supports cognitive load sharing, expert training, procedural auditing, and public communication by making uncertainty explicit, structured, and operational rather than hidden or improvised.
5 Benchmarking on the Ludwig II Case
The purpose of this section is not to provide large-scale benchmarking, which is outside the scope of the present paper, but to demonstrate that quantum abduction (QA) behaves competitively and distinctively when compared to classical abduction methods on a concrete historical case characterized by contradictory evidence and competing explanatory lines. We use the death of King Ludwig II of Bavaria (1886) as a compact abductive problem: the evidence is sparse, fragmentary, and contested, and multiple explanations remain historically defensible.
5.1 Case Setup
We consider four hypotheses:
  • H1—Suicide (self-drowning).
  • H2—Homicide (killing by guards or political actors).
  • H3—Struggle or manslaughter (accidental drowning following altercation).
  • H4—Medical episode leading to drowning.
We use seven evidence items $\mathcal{O}=\{O_{1},\dots,O_{7}\}$, expressed as natural-language snippets rather than categorical codes:
  • $O_{1}$: Autopsy reports consistent with drowning.
  • $O_{2}$: Bruising or struggle marks reported in some documents.
  • $O_{3}$: Letters and remarks interpreted as suicidal ideation.
  • $O_{4}$: Context of political removal and active opposition.
  • $O_{5}$: Witness statements are inconsistent or contradictory.
  • $O_{6}$: Prior episodes suggestive of mental health instability.
  • $O_{7}$: Conflicting timelines regarding guards’ proximity and reaction.
Hypotheses and evidence are encoded using a Sentence-BERT model [8], producing vectors $\mathbf{h}_{i}$ and $\mathbf{o}_{j}$ in a shared semantic space.
5.2 Methods Compared
We compare three approaches.
Quantum Abduction (QA)
[…]
Logic-Based Abduction (L-ABD)
We construct a small set of transparent rules:
$\{O_{3},O_{6}\}\Rightarrow H_{1},\quad\{O_{4},O_{2}\}\Rightarrow H_{2},\quad\{O_{2},O_{1}\}\Rightarrow H_{3},\quad\{O_{1}\}\Rightarrow H_{4}\;\text{(fallback)}.$
Hypotheses are scored by rule-support count, with ties broken by parsimony (least assumptions) and causal directness.
Bayesian Abduction (B-ABD)
We define a small Bayesian network with priors $p(H_{1})=0.25,p(H_{2})=0.30,p(H_{3})=0.25,p(H_{4})=0.20$ and hand-elicited likelihoods $p(O_{j}\mid H_{i})$ reflecting historically plausible support patterns. Posteriors are computed by Bayes’ rule, assuming conditional independence of evidence ($p(O_{1:m}\mid H)=\prod_{j}p(O_{j}\mid H)$).
5.3 Evaluation Protocol
We study two regimes:
R1: Full-Evidence Evaluation.
We supply all $7$ evidence items, but in random order. For each method, we compute the top-ranked hypothesis across $50$ random evidence permutations. This probes order sensitivity. Classical abduction and Bayesian inference are order-invariant; QA is not, by design, since order encodes reasoning context.
R2: Perturbation Test.
We remove $O_{2}$ (struggle indications) and repeat the experiment. This tests robustness under contradictory or missing evidence.
We report the following:
  • Top-1 distribution: proportion of runs in which each hypothesis is selected.
  • ECS (QA only): Explanatory Coherence Score $\mathrm{ECS}=\overline{e}-\big|\overline{\min(I,0)}\big|$, measuring net support vs. destructive interference.
  • Calibration proxy (QA): $\max_{i}|\alpha_{i}|^{2}$ at collapse.
  • Brier score (Bayes): posterior concentration vs. its own top prediction.
5.4 Results
The results of the benchmark are summarized in Table 5. The full benchmark implementation and reproducibility instructions are available at https://github.com/Aribertus/qa-poc (accessed on 9 December 2025). All experiments were run on a local Intel i7-8700 CPU machine with 16 GB RAM and no GPU acceleration. We report results from 50 randomized evidence orderings in two regimes: R1 (full evidence) and R2 (removal of $O_{2}$, which encodes struggle indication). Unless otherwise noted, embeddings use Sentence-BERT (all-MiniLM-L6-v2) and $\kappa_{ij}$ is initialized via sign-based exclusivity priors and smoothed by metric similarity.
Table 5: Ludwig II mini-benchmark. QA = quantum abduction (ours); L-ABD = rule-based abductive reasoning; B-ABD = Bayesian abduction. Top-1 values are fractions over 50 random evidence orders.
QA (ours) L-ABD B-ABD
Top-1 (R1) 29/50 (0.58) 0/50 (0.00) 50/50 (1.00)
Top-1 (R2) 43/50 (0.86) 0/50 (0.00) 50/50 (1.00)
ECS (QA, R1) 0.176
Conf (QA, R1) 0.374
Brier (Bayes, R1) 0.083
Interpretation.
Under full evidence (R1), QA splits between $H_{2}$ and $H_{3}$ depending on order (table reports only the $H_{2}$ Top-1 count: $29/50$), with moderate confidence ($\max_{i}|\alpha_{i}|^{2}=0.374$ on average) and positive coherence (ECS =0.176=0.176). L-ABD does not select $H_{2}$ in any run (table: $0/50$), while the Bayesian baseline (B-ABD) consistently selects $H_{2}$ ($50/50$), reflecting likelihood dominance.
Under perturbation (R2), where $O_{2}$ is removed, QA shifts toward $H_{2}$ ($43/50$ in the $H_{2}$ column), while L-ABD still does not select $H_{2}$ and B-ABD remains at $H_{2}$ ($50/50$). Overall, QA manages uncertainty gracefully (maintaining superposition until constraint accumulates), whereas the classical baselines either hard-select or revert to priors.
5.5 Summary
This compact benchmark shows that quantum abduction is as follows:
  • Competitive with classical abductive methods under fully informative conditions;
  • More robust when evidence is contradictory or incomplete;
  • Uniquely capable of modeling order/context effects and hybrid explanatory outcomes.
A full-scale benchmark suite (multiple domains, larger hypothesis sets, robustness to noise, expert-elicited interference structures, and human evaluation of hybrid explanations) is left for future work, as discussed in Section 7.
@@ 6.3 Set-Covering Approaches
Set-covering abduction, pioneered by Reggia and colleagues [21, 22, 23], frames explanation as a combinatorial optimization task: find minimal sets of causes that “cover” observed effects. Widely applied in medical diagnosis and fault detection, this approach is computationally tractable and intuitively appealing in domains with clear causal mappings. Variants such as Parsimonious Covering Theory further emphasized minimality as a criterion of explanatory adequacy.
@@ 6.4 Integrative Assessment
Across these traditions, abduction is implemented in structurally different ways. To enable their direct comparison as in Table 6, we assess them along four shared dimensions:
  • Hypothesis representation: whether hypotheses are treated as discrete alternatives or as jointly coexisting explanatory states.
  • Conflict handling: whether contradictions are eliminated, suspended, or allowed to interact.
  • Update dynamics: how new evidence changes explanatory states.
  • Outcome structure: whether reasoning terminates in selection, ranking, or synthesis.
Table 6: Unified comparison across abductive paradigms. Classical approaches select among hypotheses, while quantum abduction allows coexistence and interaction prior to collapse.
Dimension Logic-Based Bayesian Set-Covering Quantum Abduction (Ours)
Hypothesis representation Discrete alternatives Weighted alternatives Minimal covering sets Coexisting amplitudes (superposition)
Conflict handling Eliminated by consistency Resolved by posterior dominance Removed via minimality constraints Retained with constructive/destructive interference
Update dynamics Rule application and pruning Likelihood reweighting Constraint-/cost-based elimination Amplitude rotation and phase interaction
Outcome structure Single minimal explanation Most probable explanation (MPE) Parsimonious cover Collapse to dominant or hybrid explanation
6.5 Alternative and Hybrid Approaches
Beyond the three main paradigms reviewed above—logic-based, Bayesian, and set-covering—other research lines approximate abductive reasoning through mechanisms that relax bivalence, incorporate argumentative dynamics, or combine neural and symbolic representations. Although these approaches introduce valuable flexibility, they remain limited in their ability to model superposition, interference, and synthesis within a unified formal framework.
Fuzzy Logic Systems.
Fuzzy logic provides a graded alternative to classical deduction by replacing Boolean truth values with degrees of membership [24, 25]. In abductive contexts, fuzzy inference allows for partial matching between observations and hypotheses, accommodating uncertainty in linguistic and perceptual descriptions. However, fuzzy systems operate on scalar degrees of truth and lack the vectorial and phase-sensitive interactions that characterize quantum interference. As a result, while fuzzy logic captures vagueness, quantum abduction captures the interaction among competing explanatory tendencies.
Argumentation Frameworks.
Another major non-classical approach involves formal argumentation and defeasible reasoning. Dung’s abstract argumentation framework [26] and its numerous descendants model the interplay of conflicting arguments through attack and defense relations, enabling structured deliberation and justification. These systems effectively represent contradiction and defeat but typically resolve conflict through the selection of stable or preferred extensions. Quantum abduction, by contrast, maintains superposed explanatory states in which conflicting hypotheses may coexist and influence one another until contextual constraints induce collapse, thus supporting a non-eliminative form of reasoning.
Neural-Symbolic Integration.
A third trajectory aims to bridge symbolic and sub-symbolic inference through neural-symbolic computation [27, 28]. These frameworks embed logical rules within neural architectures, enabling differentiable reasoning under uncertainty. While this integration has proven powerful for knowledge representation and learning, it remains primarily a fusion architecture: logic guides neural adaptation, but the symbolic layer itself does not exhibit quantum-like contextuality or interference. Quantum abduction complements neural-symbolic integration by providing a principled mathematical model—rooted in Hilbert space semantics—for managing contradictory symbolic information within continuous neural embeddings.
Summary.
In summary, these alternative approaches enrich the abductive landscape by addressing uncertainty, conflict, and hybrid reasoning. Yet none explicitly capture the non-classical compositionality central to human explanatory cognition—the coexistence and dynamic blending of partial hypotheses. Quantum abduction provides a coherent formalism that encompasses these features within a single, interference-enabled representational framework.
7 Conclusions, Outlook, and Future Work
This paper has presented quantum abduction as a coherent extension of abductive reasoning beyond the eliminative logics that dominate classical, Bayesian, and set-covering paradigms. Where those traditions converge on ranking and discarding hypotheses, quantum abduction preserves them in a dynamic superposition that allows constructive and destructive interference, culminating in collapse only when contextual evidence supports a coherent, possibly hybrid, synthesis.
The framework thus reinterprets explanation as a process of synthesis through interaction rather than selection through exclusion. It accounts for how human reasoning sustains ambiguity, navigates contradiction, and gradually achieves insight through the tension of partially competing interpretations. At the computational level, this paradigm integrates semantic embeddings, projection operators, and interference matrices into an amplitude-based dynamics capable of representing explanatory coexistence.
Conceptually and architecturally, this work stands as a step within a broader research program that seeks to overcome eliminativism in reasoning—across logic, cognition, and artificial intelligence. Quantum abduction offers not a speculative draft but an operational framework that is already formalized at the representational level and computationally implemented in proof-of-concept systems. Its orientation is centaurian [10, 9]: human experts remain within the loop, guiding semantic alignment and interpretive synthesis, while the computational substrate maintains the full superpositional structure inaccessible to unaided intuition. In this sense, the framework realizes a concrete model of human–AI complementarity, aligning with emerging paradigms of agentic and hybrid intelligence.
7.1 Current Scope and Limitations
The evaluation presented here is illustrative rather than exhaustive. Benchmarks such as the Ludwig II study demonstrate the feasibility of the approach but do not yet provide statistical generalization. Interference coefficients $\kappa_{ij}$ are set heuristically; large-scale learning from empirical data remains to be achieved. Likewise, while prototype implementations exist, no controlled human studies have yet quantified improvements in interpretive accuracy or decision quality. Finally, computational cost grows with the dimensionality of the hypothesis space, motivating research into sparse and approximate formulations.
7.2 Future Directions
Future work will extend the current framework along the following three convergent axes:
  1. 1.
    Formal and Logical Development. We will elaborate the proof-theoretic and model-theoretic foundations of quantum abduction, establishing soundness and completeness properties for amplitude-based reasoning. This involves importing techniques from quantum logic and category theory to represent composition, interference, and collapse within a unified semantic calculus.
  2. 2.
    Computational Expansion. A scalable software library and benchmark suite will be released as part of the ongoing open-source program. Optimizing $\kappa_{ij}$ through learning, expanding to higher-dimensional embeddings, and integrating retrieval-augmented generation (RAG) pipelines will support domain adaptation across forensic, clinical, and scientific settings.
  3. 3.
    Human-in-the-Loop Validation. Controlled studies with experts will test how the quantum abductive assistant influences reasoning transparency, confidence calibration, and interpretive synthesis. These experiments will advance hybrid decision frameworks where collapse is a collaborative, not unilateral, outcome.
7.3 Programmatic Outlook
The quantum abductive approach contributes to the ongoing effort to construct reasoning systems that do not merely simulate deductive or statistical inference, but actively support the generative and exploratory dimensions of human thought. Its integration of symbolic clarity, probabilistic nuance, and superpositional synthesis marks a shift from the competition of explanations to their entanglement. In this sense, quantum abduction represents both a conceptual unification and a methodological bridge—one that connects the epistemic pluralism of human inquiry with the algorithmic precision of modern AI.
Funding:
@@ 7.3 Programmatic Outlook
Data Availability Statement:
The code and data for the proof-of-concept implementation (Ludwig II benchmark) described in this manuscript are available in the public repository https://github.com/Aribertus/qa-poc (accessed on 9 December 2025).
Conflicts of Interest:
The authordeclares no conflicts of interest.The funders had no role in the design of the study; in the collection, analyses, or interpretation of data; in the writing of the manuscript; or in the decision to publish the results.
Acknowledgments
The author is grateful to Hervé Gallaire and the reviewers for comments and suggestions which have been of great help in improving this article. While preparing this manuscript, the author used ChatGpt 5 and DeepSeek to guide him in LaTeX formatting, particularly in using the TikZ package for creating graphs, as well as in the polishing of the text, aiming to make it as fluent and communicable as possible. The author reviewed and edited the resulting output and takes full responsibility for the content of this publication.
Appendix A Glossary of Key Terms
Table 7: Key terms and concepts in quantum abduction
Term Definition
Transformer A neural architecture for encoding sequences into context-sensitive vector representations. Used here to embed hypotheses and observations in a shared semantic space.
LLM (Large Language Model) A generative model producing natural-language text. In quantum abduction, used only for articulating hybrid hypotheses, not for scoring or selecting them.
Superposition The coexistence of multiple hypotheses as weighted amplitudes in the explanatory state, rather than treating them as mutually exclusive alternatives.
Interference The phenomenon whereby hypotheses can reinforce (constructive) or diminish (destructive) each other’s explanatory power through semantic interaction.
Amplitude ($\alpha_{i}$) The complex coefficient associated with hypothesis $H_{i}$ in the superposition, with $|\alpha_{i}|^{2}$ representing its relative weight.
Collapse The convergence of the superposed state to a dominant explanation or synthesized hybrid when coherence exceeds a threshold.
Projection The semantic alignment between an observation and a hypothesis, computed as cosine similarity in the embedding space.
Entanglement In our context (distinct from physics), the semantic interdependence between hypotheses where their meanings and explanatory power shift based on co-activation.
Mix Operator The mathematical function that combines high-amplitude hypotheses into hybrid explanations through weighted semantic blending.
Coherence A measure of how concentrated the explanatory state is, typically $C(\Psi)=\max_{i}|\alpha_{i}|^{2}$.
Epistemic vs. Ontological Our superposition is epistemic (about our knowledge/uncertainty), not ontological (about reality creating multiple worlds).
$\kappa_{ij}$ Interference coefficient encoding domain-specific interaction strength between hypotheses $i$ and $j$.
Appendix B Formal Framework and Derivations
B.1 State, Projection, and Interference
State Representation
[…]
Semantic Embedding
$\mathbf{h}_{i},\mathbf{o}_{j}\in\mathbb{R}^{d},\qquad\mathbf{h}_{i}=\text{SBERT}(\text{text}(H_{i})),\;\;\mathbf{o}_{j}=\text{SBERT}(\text{text}(O_{j}))$ (4)
Projection Operator
$\langle H_{i}|O_{j}\rangle=\cos(\mathbf{h}_{i},\mathbf{o}_{j})=\frac{\mathbf{h}_{i}\cdot\mathbf{o}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{o}_{j}\|}$ (5)
Interference Matrix
$I_{ij}=\kappa_{ij}\cdot\text{sim}(H_{i},H_{j}),\qquad\kappa_{ij}\in[-1,1]$ (6)
B.2 Amplitude Dynamics and Coherence
Update Rule and Normalization
$\tilde{\alpha}_{i}^{(t+1)}=\alpha_{i}^{(t)}+\eta\left[e_{i}^{(t)}+\sum_{k\neq i}I_{ik}\alpha_{k}^{(t)}\right],\qquad\alpha^{(t+1)}=\frac{\tilde{\alpha}^{(t+1)}}{\|\tilde{\alpha}^{(t+1)}\|_{2}}$ (7)
Coherence and Collapse
[…]
B.3 Mix Operator and Synthesis
Mix Operator
$\text{mix}(H_{i},H_{j})=\lambda_{ij}\mathbf{h}_{i}+(1-\lambda_{ij})\mathbf{h}_{j},\qquad\lambda_{ij}=\frac{|\alpha_{i}|^{2}}{|\alpha_{i}|^{2}+|\alpha_{j}|^{2}}$ (9)
For natural language articulation, we utilize template-based synthesis for simple merges and LLM-based critique to refine complex hybrids.
B.4 Implementation Sketch and Complexity
Reference Sketch (pseudo-code).
Input: alpha[1..n], I[1..n,1..n], evidence e[1..n], eta in (0,1]
for i in 1..n:
[…]
    alpha[k] )
Normalize: alpha = alpha_tilde / ||alpha_tilde||_2
Collapse if max_i |alpha[i]|ˆ2 > tau; else continue
Estimating $I_{ij}$.
Initialize from cosine similarity; set $\kappa_{ij}$ via domain priors, weak supervision (optimize ECS/HQI on solved cases), or expert elicitation.
Complexity.
Per step $O(n^{2})$ for interference + $O(n)$ for projection aggregation; with $m$ observations, $O(n^{2}m)$.
Appendix C Computational Sketch and Implementation Notes
C.1 State, Evidence, Interference
Let $\mathcal{H}=\{H_{i}\}_{i=1}^{n}$ and $\mathcal{O}=\{O_{j}\}_{j=1}^{m}$. Maintain
[…]
Embed hypotheses and observations with a sentence encoder (e.g., Sentence-BERT [8]) to obtain vectors $\mathbf{h}_{i},\mathbf{o}_{j}\in\mathbb{R}^{d}$. Use cosine similarity as a proxy for $\langle H_{i}\mid O_{j}\rangle$.
Evidence projection for a new batch is as follows $\mathcal{O}^{(t)}$:
$e_{i}^{(t)}\;=\;\max_{O_{j}\in\mathcal{O}^{(t)}}\frac{\mathbf{h}_{i}\cdot\mathbf{o}_{j}}{\|\mathbf{h}_{i}\|\,\|\mathbf{o}_{j}\|}.$
Interference matrix $I\in\mathbb{R}^{n\times n}$ models hypothesis interaction. A simple, effective choice is the following:
$I_{ii}=1,\qquad I_{ij}=\mathrm{cos\_sim}(\mathbf{h}_{i},\mathbf{h}_{j})\cdot\kappa_{ij},\;\;i\neq j,$
with $\kappa_{ij}\in[-1,1]$ permitting constructive (>0>0) or destructive (<0<0) coupling. In data-poor settings set $\kappa_{ij}=1$ (pure similarity). When expert priors are available, hand-tune signs to encode known exclusivities or complementarities.
C.2 Amplitude Update and Collapse
With learning rate $\eta\in(0,1)$
$\tilde{\alpha}_{i}^{(t+1)}\;=\;\alpha_{i}^{(t)}+\eta\Big(e_{i}^{(t)}+\sum_{k\neq i}I_{ik}\,\alpha_{k}^{(t)}\Big),\qquad\alpha_{i}^{(t+1)}\;=\;\frac{\tilde{\alpha}_{i}^{(t+1)}}{\sqrt{\sum_{\ell}|\tilde{\alpha}_{\ell}^{(t+1)}|^{2}}}.$
Define a coherence functional that grows when one (or a consistent subset) dominates:
$\mathcal{C}^{(t)}\;=\;\max_{i}|\alpha_{i}^{(t)}|^{2}\quad\text{or}\quad\mathcal{C}^{(t)}\;=\;\sum_{i}|\alpha_{i}^{(t)}|^{4}.$
Collapse when $\mathcal{C}^{(t)}\geq\tau$ or when a decision deadline arrives. If several $|\alpha_{i}|$ remain high and $I_{ij}>0$, return a hybrid synthesis $H_{i\wedge j}$.
C.3 Synthesis operator (hybrids)
Given two high-mass hypotheses $H_{i},H_{j}$, form a content-level blend
$\mathrm{mix}(\mathbf{h}_{i},\mathbf{h}_{j})=\lambda\,\mathbf{h}_{i}+(1-\lambda)\,\mathbf{h}_{j},\quad\lambda=\frac{|\alpha_{i}|^{2}}{|\alpha_{i}|^{2}+|\alpha_{j}|^{2}},$
and instantiate a natural-language synthesis using a controlled LLM prompt that lists (i) active hypotheses and amplitudes, (ii) salient observations and projections, (iii) entries of $I_{ij}$ that justify merger or exclusion. This yields faithful, auditable text (see [11] for a clinical instance and [2] for a strategic instance).
C.4 Implementation Note
A lightweight pipeline suffices the following:
  1. 1.
    Sentence-BERT embeddings for $\mathbf{h}_{i},\mathbf{o}_{j}$ (384–768d).
  2. 2.
    Evidence scores $e_{i}$ via cosine similarity.
  3. 3.
    $I_{ij}$ from similarity; optionally sign $\kappa_{ij}$ with expert priors (exclusivity vs. complementarity).
  4. 4.
    Iterate the update; monitor $\mathcal{C}^{(t)}$; stop at $\tau$ or deadline.
  5. 5.
    If hybrid, call the synthesis operator with $(\alpha,I,e)$ as structured context.